# Optimal Low-Depth Quantum Signal-Processing Phase Estimation

### v1

@@ Abstract
Quantum effects like entanglement and coherent amplification can be used to drastically enhance the accuracy of quantum parameter estimation beyond classical limits. However, challenges such as decoherence and time-dependent errors hinder Heisenberg-limited amplification. We introduce Quantum Signal-Processing Phase Estimation algorithms that are robust against these challenges and achieve optimal performance as dictated by the Cramér-Rao bound. These algorithms use quantum signal transformation to decouple interdependent phase parameters into largely orthogonal ones, ensuring that time-dependent errors in one do not compromise the accuracy of learning the other. Combining provably optimal classical estimation with near-optimal quantum circuit design, our approach achieves an unprecedented standard deviation accuracy of $10^{-4}$ radians for estimating unwanted swap angles in superconducting two-qubit experiments, using low-depth ($<10$) circuits. This represents up to two orders of magnitude improvement over existing methods. Theoretically and numerically, we demonstrate the optimality of our algorithm against time-dependent phase errors, observing that the variance of the time-sensitive parameter $\varphi$ scales faster than the asymptotic Heisenberg scaling in the small-depth regime. Our results are rigorously validated against the quantum Fisher information, confirming our protocol’s ability to achieve unmatched precision for two-qubit gate learning.
Quantum metrology’s efficiency is fundamentally influenced by two critical factors: the Heisenberg limit, which defines how accuracy scales with quantum resources, and the coefficients of this scaling. While a variety of quantum metrology strategies [14, 23, 2] successfully adhere to the Heisenberg scaling, the real challenge lies in achieving or even addressing optimality in the scaling coefficients with realistic constraints. This aspect is particularly vital for applications in quantum error correction, where achieving fault-tolerant thresholds demands exceptionally high accuracy in quantum gate characterization. The necessity for deep circuitry, a significant hurdle in practical applications, stems directly from the lack of optimality in these scaling coefficients. This inefficiency is compounded by the challenges of finite coherence times and the amplification of drift errors from low-frequency noise or control fluctuations. Therefore, current quantum metrology protocols, limited to accuracy levels between $10^{-2}$ and $10^{-3}$ radians for estimating gate angles, often fall short of the accuracy ($\sim 10^{-4}$) needed to verify the crossing of fault-tolerant error threshold for quantum error correction and other near-term quantum applications.
@@ Results
We propose a metrology protocol that is, by design, robust against realistic time-dependent errors and only requires shallow ($<10$) circuits to achieve up to two orders of magnitudes of improvement over existing methods in the precision of gate-parameter estimates. We harness the analytical structure of a class of quantum-metrology circuits using a theoretical toolbox from classical signal processing [12, 30], Quantum Signal Processing (QSP) [17, 10, 33] and polynomial analysis [21]. QSP allows us to treat the inherent quantum dynamics as input quantum signals and perform universal transformations on the input to realize targeted quantum dynamics as output. Classical signal processing provides methods analyzing these transformed signals to produce robust estimations. We propose a general gate model, which we term $U$-gate model, that encapsulates two-level invariant subspace structure in the native gate sets of superconducting, neutral atoms, and ion trap quantum computers. We parameterize the subspace of interest in our model $U$-gates with a set of angle parameters, and provide a metrology algorithm on learning the swap angle $\theta$ and the phase difference $\varphi$.
@@
Our metrology algorithm, which we term Quantum Signal-Processing Phase Estimation (QSPE), separates the estimation of the parameter-free from time-dependent errors ($\theta$) from that which is affected by time-dependent drift ($\varphi$). Interestingly, the parameter $\varphi$ variance shrinks faster than Heisenberg scaling concerning circuit depth in the pre-asymptotic low-depth regime of experimental interest. We analyze the stability of our protocol in the presence of realistic experimental noise and sampling errors. We prove that our method achieves the Cramér-Rao lower bound in the presence of sampling errors and achieves up to $10^{-4}$ STD accuracy in learning swap angle $\theta$ in both simulation and experimental deployments on superconducting qubits. To the best of our knowledge, we provide the first evaluation of a metrology protocol’s quantum Fisher information (QFI) and show that our approach is a factor of two above the quantum CRLB (QCRLB). Furthermore, we demonstrate an interesting transition of the optimal metrology variance scaling as a function of circuit depth $d$ from the pre-asymptotic regime $d\ll 1/\theta$ to the Heisenberg limit $d\to\infty$.
@@
We summarize the main results of our metrology algorithm and start by defining the metrology problem, the learning of a general $U$-gate, followed by an analysis of the QSP circuit with $U$-gates used in our algorithm. Building upon these closed-form results, we propose a phase estimation method combining Fourier analysis with QSP to separate the two gate parameters of interest in their functional forms. Our new estimation algorithm enables fast and deterministic data post-processing using only direct linear algebra operations rather than iterative black-box optimizations used in multi-parameter robust phase estimation [14, 23]. Furthermore, the separation of inference of $\theta$ and $\varphi$ improves the robustness of the phase estimation method against dominantly time-dependent error on one of the gate parameters. The analysis and modeling of Monte Carlo sampling error also indicate that our phase estimation method achieves the fundamental quantum metrology optimality in a practical regime against realistic errors for near-term devices. We also provide a comprehensive mathematical analysis of methods based on robust phase estimation [2, 23], and prove that the vulnerability of phase angle $\varphi$ to time-dependent errors ultimately renders the estimation accuracy of the swap angle $\theta$ exponentially worse than Heisenberg limit. In addition, we proposed a noise-robust QSPE protocol that enables the estimation of gate parameters even when the gate angles fall outside the confidence regime for phase estimation. Lastly, we demonstrated QSPE experimentally on 34 superconducting qubits using the Google Quantum AI team’s hardware, showcasing up to two orders-of-magnitude improvement over previous methods, achieving $10^{-4}$ two-qubit phase estimation accuracy in practice. Lastly, we include an empirical noise-robust QSPE protocol that enables the estimation of gate parameters even when the gate angles fall outside of the approximation regime for phase estimation.
@@
Previous metrology methods [14, 23, 2] based on optimal measurements [5] for achieving the Heisenberg limit fall short of providing sufficient accuracy in $\theta$ when $\theta\ll 1$. Two significant factors limit these traditionally-regarded “optimal” metrology schemes. First, the accuracy in $\theta$ depends on the amplification factor, i.e., maximum circuit depth. The relatively low qubit coherence times of superconducting qubits render randomization-based quantum gate learning techniques [15, 19, 3] impractical due to their inefficient circuit depths. Techniques based on robust phase estimation can require prohibitive depths to achieve a meaning full signal-to-noise ratio for small $\theta$ and require iterative black-box optimizations for their estimators [23] instead of fast, deterministic post-processing for single-parameter phase estimation [14]. Second, time-dependent unitary error in $\varphi$ is prevalent in architectures like Google’s superconducting quantum computers [27], which invalidates basic assumptions in traditionally optimal and Heisenberg-limit-achieving metrology schemes.
@@
In this work, we provide a low-depth phase estimation method for estimating the angles in some invariant subspace of an unknown $U$-gate when the swap angle is extremely small, of order below $10^{-3}$, while facing realistic time-dependent phase errors in $\varphi$. The phase estimation method leverages the structure of periodic circuits analyzed by classical and quantum signal processing and provides a framework to engineer quantum metrology from the perspective of universal quantum signal transformation. We call this new type of metrology method Quantum Signal-Processing Phase Estimation (QSPE). Let $\omega$ be a tunable phase parameter and {|$\{\ket{0_{\ell}},\ket{1_{\ell}}\}$ be the logical basis of the two-level space of interest. Then, QSPE measures the transition probabilities of the quantum circuits corresponding to logical Bell states |$\ket{+_{\ell}}:=\frac{1}{\sqrt{2}}\left(\ket{0_{\ell}}+\ket{1_{\ell}}\right)$ and |$\ket{\mathrm{i}_{\ell}}:=\frac{1}{\sqrt{2}}\left(\ket{0_{\ell}}+\mathrm{i}\ket{1_{\ell}}\right)$. The transition is measured with respect to the logical basis state |$\ket{0_{\ell}}$. We depict the quantum circuit for QSPE in Figure 1 with an exemplified two-qubit $U$-gate for simplicity. In the example, the two-level subspace is set to the single-excitation subspace with basis |$\ket{0_{\ell}}=\ket{01}$ and |$\ket{1_{\ell}}=\ket{10}$. Then, the logical Bell state coincides with the conventional Bell state. We remark that the quantum circuit for QSPE can be generalized to multi-qubit cases following the recipe outlined in this paragraph. Details and the quantum circuit for QSPE in a general setup are provided in Appendix B. The transition probability corresponding to the logical Bell state $\ket{+_{\ell}}$ is denoted as $p_{X}(\omega;\theta,\varphi,\chi)$, and that corresponding to the logical Bell state |$\ket{\mathrm{i}_{\ell}}$ is denoted as $p_{Y}(\omega;\theta,\varphi,\chi)$.
Figure 1: Quantum circuit for QSPE with an exemplified two-qubit $U$-gate. The input quantum state is prepared to be Bell state in either $\ket{+_{\ell}}$ or |$\ket{\mathrm{i}_{\ell}}$ according to the type of experiment. The quantum circuit enjoys a periodic structure of the unknown $U$-gate and a tunable $Z$ rotation.
@@
As outlined in Appendix C, the reconstructed function derived from measurement probabilities admits an approximated expansion $\mathfrak{h}(\omega;\theta,\varphi,\chi)=\sum_{-d+1}^{d-1}c_{k}e^{2\mathrm{i}k\omega}$ and when $d\theta\leq 1$, the coefficients are
@@
Due to Fourier expansion, sampling the reconstructed function on $(2d-1)$ distinct $\omega$-points is sufficient to characterize its information completely. The second important consequence of this result is that the dependencies on $\theta$ and $\varphi$ are completely separated into the amplitude and the phase of the Fourier coefficients, respectively. The estimation problems of $\theta$ and $\varphi$ are then reduced to two independent linear regression problems. As $\chi$ is not considered, we apply a sequential phase difference to distill the angle $\varphi$:
$\vec{\Delta}=(\Delta_{0},\cdots,\Delta_{d-2})^{\top},\Delta_{k}:=\mathsf{phase}(c_{k}\overline{c_{k+1}})=2\varphi.$ (8)
Considering the Monte Carlo sampling error due to the finite number of measurements, we derive in Appendix D the linear-regression-based estimators of the relevant angles:
$\hat{\theta}=\frac{1}{d}\sum_{k=0}^{d-1}\absolutevalue{c_{k}}\quad\text{ and }\quad\hat{\varphi}=\frac{1}{2}\frac{\vec{\mymathbb{1}}^{\top}\mathfrak{D}^{-1}\vec{\Delta}}{\vec{\mymathbb{1}}^{\top}\mathfrak{D}^{-1}\vec{\mymathbb{1}}}.$ (9)
Here, $\vec{\mymathbb{1}}$ is an all-one vector and $\mathfrak{D}$ is a $(d-1)$-by-$(d-1)$ constant tridiagonal matrix which is coincides with the discrete Laplacian of a central finite difference form (see Definition 17 in Appendix D for more details). The structure of $\mathfrak{D}$ comes from differentiating experimental noises when applying sequential phase difference. To obtain $\hat{\chi}$, we defer the task to the metrology circuit in [2, Fig. S5] and do not use QSPE for the task. The main workflow of the QSPE is depicted in Figures 1 and 2.
{algorithm}
[htbp] Inferring unknown angles in $U$-gate with extremely small swap angle using QSPE
  Initiate a complex-valued data vector $\vec{\mathfrak{h}}\in\mathbb{C}^{2d-1}$.
  for $j=0,1,\cdots,2d-2$ do
     Set the tunable $Z$-phase angle as $\omega_{j}=\frac{j}{2d-1}\pi$.
     Set $\vec{\mathfrak{h}}_{j}\leftarrow p_{X}(\omega_{j})-\frac{1}{2}+\mathrm{i}\left(p_{Y}(\omega_{j})-\frac{1}{2}\right)$.
  end for
  Compute the Fourier coefficients $\vec{c}=\mathsf{FFT}\left(\vec{\mathfrak{h}}\right)$.
  Output: Estimators $\hat{\theta},\hat{\varphi}$.
Figure 2: Flowchart of main procedures in QSPE.
@@
Their biases and variances measure the performance of the statistical estimators. In Section D.2, we derive the performance of QSPE estimators with the following theorem by treating QSPE as linear statistical models. Furthermore, in Appendix F, we show that QSPE estimators in Equation 9 are optimal by saturating the Cramér-Rao lower bound (CRLB) of the estimation problem.
@@
According to the framework developed in Ref. [11], the variance of any quantum metrology is lower bound by the Heisenberg limit. It indicates that in our experimental setup when $d$ is large enough, optimal variance scales as $1/(d^{3}M)$. This seemingly contradicts Theorem 2, where the variance of QSPE $\varphi$-estimator can achieve $1/(d^{4}M)$. This counterintuitive conclusion is due to the pre-asymptotic regime $d\ll 1/\theta$. In Appendix F, we analyze the CRLB of QSPE. The optimal variance given by CRLB is exactly solvable in the pre-asymptotic regime $d\ll 1/\theta$.
The key reason behind such faster than Heisenberg limit scaling in pre-asymptotic regime depends on the unique structure of the QSPE circuit: the measurement outcome (Equation 19) concentrates around a constant value regardless of the gate parameter values. Yet when $d$ is large enough to pass to the asymptotic regime, measurement probabilities will take arbitrary values. Furthermore, the analysis of the CRLB suggests that the optimal asymptotic variance agrees with the Heisenberg limit when $d$ is alrge enough. This nontrivial transition of optimal variance is theoretically analyzed and numerically justified in Appendix F. We summarize this nontrivial transition of the optimal variance scaling of QSPE as a phase diagram in Figure 3(a). To numerically demonstrate the transition, we compute the exact CRLB of QSPE when $\theta=1\times 10^{-2}$ and $\theta=1\times 10^{-3}$. In Figure 3(b), the slope of the curve in log-log scale exhibits a clear transition before and after $d=1/\theta$, which supports the phase diagram in Figure 3(a). Detailed theoretical and numerical discussions of the transition are carried out in Appendix F.
Figure 3: A nontrivial transition of the optimal variance in solving QSPE. The theoretical analysis of the transition is in Appendix F. (a) Phase diagram showing the nontrivial transition of the optimal variance in solving QSPE. QSPE estimators attain the optimal variance in the pre-asymptotic regime. (b) Cramér-Rao lower bound (CRLB) and the theoretically derived optimal variance in the pre-asymptotic regime. The single-qubit phases are set to $\varphi=\pi/16$ and $\chi=5\pi/32$. The number of measurement samples is set to $M=1\times 10^{5}$.
@@
Analyzing Cramér-Rao bounds suggests the optimality of a quantum metrology protocol or the suboptimality leading to further improvement. Given an initialization and measurement, the optimality lies in the analysis of the classical CRLB, which investigates the most information one can retrieve from measurement probabilities. As outlined in Section F.1, the CRLBs are solvable when $d\ll 1/\theta$, which exactly agree with the variance of our estimators derived in Theorem 2. The optimality of our estimator is also validated from numerical simulation depicted in Figure 3(b). Such agreement reveals the optimality of our data post-processing. Although linear-regression-based estimators are used, this linearization does not sacrifice the information retrieval in the experimental data. Furthermore, in contrast to other iterative inference methods, our estimators directly estimate angles using basic linear algebra operations, to which stability and fast processing are credited.
Despite the informative indication by analyzing CRLB, it cannot provide direct suggestions on improving initialization and measurement. Such generalization demands the switch to quantum Cramér-Rao lower bound (QCRLB) which requires upper bounding the quantum Fisher information (QFI). As a quantum analog of the classical Fisher information, QFI lies in the center of quantum metrology by providing a fundamental lower bound on the accuracy one can infer from the system of a given resource limit. According to the analysis in [16], the QFI is an upper bound on the Fisher information over all possible measurements. For brevity, we only consider the inference of $\theta$ and hold all other unknown parameters constant in the analysis. However, our analysis can be generalized to the multiple parameter inference by adopting the multi-variable QFI in Ref. [16]. In Section F.4, we derive that the average QFI is upper bounded by an integral:
@@
Here, the integrand gauges the information contained in an experiment with angle $\omega$. The integral stands for the use of equally spaced $\omega$ samples due to the absence of accurate information of $\varphi$. It is worth noting that the integrand is sharply peaked at $4d^{2}$ when $\omega=\varphi$ which is also referred to as the phase-matching condition. The missing information of $\varphi$ downgrades the average QFI from $4d^{2}$ to $4d$. However, as discussed in the following subsection, the lack of information about $\varphi$ in existing methods can potentially significantly degrade the Fisher information to $\mathcal{O}(\log(d))$, thereby severely impeding the achievement of Heisenberg-limit scaling in estimation accuracy. Furthermore, this also suggests a finer estimation of $\theta$ when some rough information of $\varphi$ is provided either as a priori or from some preliminary estimation. This improvement is discussed in Section D.3. Consequently, the QCRLB of the QSPE formalism is
@@
Compared with Theorem 2, we see differentiation in a constant suboptimal factor of $2$, which is explainable. Note that we use two logical Bell states to perform experiments. The advantage is the experimental probabilities of these two experiments form a conjugate pair to reconstruct a complex function that for ease of analysis. This complex function and its properties (see Theorems 9 and 11) eventually lead to a simple robust statistical estimator requiring only light computation. In contrast, the data generated from the initialization of one Bell state still contains full information on the parameters to be estimated. However, the highly nonlinear and oscillatory dependency renders the practical inference challenging. Hence, the factor of $2$ is due to the use of a pair of Bell states. Although the QFI indicates that inference variance can be lower by removing such redundancy in the initialization, the nature of ignoring practical ease makes it hard to achieve.

### v2

@@ Abstract
Quantum effects like entanglement and coherent amplification can be used to drastically enhance the accuracy of quantum parameter estimation beyond classical limits. However, challenges such as decoherence and time-dependent errors hinder Heisenberg-limited amplification. We introduce Quantum Signal-Processing Phase Estimation algorithms that are robust against these challenges and achieve optimal performance as dictated by the Cramér-Rao bound. These algorithms use quantum signal transformation to decouple interdependent phase parameters into largely orthogonal ones, ensuring that time-dependent errors in one do not compromise the accuracy of learning the other. Combining provably optimal classical estimation with near-optimal quantum circuit design, our approach achieves a standard deviation accuracy of $10^{-4}$ radians for estimating unwanted swap angles in superconducting two-qubit experiments, using low-depth ($<10$) circuits. This represents up to two orders of magnitude improvement over existing methods. Theoretically and numerically, we demonstrate the optimality of our algorithm against time-dependent phase errors, observing that the variance of the time-sensitive parameter $\varphi$ scales faster than the asymptotic Heisenberg scaling in the small-depth regime. Our results are rigorously validated against the quantum Fisher information, confirming our protocol’s ability to achieve unmatched precision for two-qubit gate learning.
Quantum metrology’s efficiency is fundamentally influenced by two critical factors: the Heisenberg limit, which defines how accuracy scales with quantum resources, and the coefficients of this scaling. While a variety of quantum metrology strategies [16, 26, 3] successfully adhere to the Heisenberg scaling, the real challenge lies in achieving or even addressing optimality in the scaling coefficients with realistic constraints. This aspect is particularly vital for applications in quantum error correction, where achieving fault-tolerant thresholds demands exceptionally high accuracy in quantum gate characterization. The necessity for deep circuitry, a significant hurdle in practical applications, stems directly from the lack of optimality in these scaling coefficients. This inefficiency is compounded by the challenges of finite coherence times and the amplification of drift errors from low-frequency noise or control fluctuations. Therefore, current quantum metrology protocols, limited to accuracy levels between $10^{-2}$ and $10^{-3}$ radians [26, 3] for estimating gate angles, often fall short of the accuracy ($\sim 10^{-4}$) needed to verify the crossing of fault-tolerant error threshold for quantum error correction and other near-term quantum applications.
@@
In this work, we propose a metrology protocol that is, by design, robust against realistic time-dependent errors and only requires shallow ($<10$) circuits to achieve up to two orders of magnitudes of improvement over existing methods in the precision of gate-parameter estimates.
@@ Results
We harness the analytical structure of a class of quantum-metrology circuits using a theoretical toolbox from classical signal processing [14, 33], Quantum Signal Processing (QSP) [19, 11, 37] and polynomial analysis [23]. QSP allows us to treat the inherent quantum dynamics as input quantum signals and perform universal transformations on the input to realize targeted quantum dynamics as output. Classical signal processing provides methods for analyzing these transformed signals to produce robust estimations. We propose a general gate model, which we term $U$-gate model, that encapsulates two-level invariant subspace structure in the native gate sets of superconducting, neutral atoms, and ion trap quantum computers. We parameterize the subspace of interest in our model $U$-gates with a set of angle parameters, and provide a metrology algorithm for learning the swap angle $\theta$ and the phase difference $\varphi$.
@@
Our metrology algorithm, which we term Quantum Signal-Processing Phase Estimation (QSPE), separates the estimation of the parameter-free from time-dependent errors ($\theta$) from that which is affected by time-dependent drift ($\varphi$). Interestingly, the parameter $\varphi$ variance shrinks faster than Heisenberg scaling concerning circuit depth in the pre-asymptotic low-depth regime of experimental interest. We analyze the stability of our protocol in the presence of realistic experimental noise and sampling errors. We prove that our method achieves the Cramér-Rao lower bound in the presence of sampling errors and achieves up to $10^{-4}$ STD accuracy in learning swap angle $\theta$ in both simulation and experimental deployments on superconducting qubits. We provide the evaluation of metrology protocol’s quantum Fisher information (QFI) and show that our approach is a factor of two above the quantum CRLB (QCRLB). Furthermore, we demonstrate an interesting transition of the optimal metrology variance scaling as a function of circuit depth $d$ from the pre-asymptotic regime $d\ll 1/\theta$ to the Heisenberg limit $d\to\infty$.
@@
We summarize the main results of our metrology algorithm and start by defining the metrology problem, the learning of a general $U$-gate, followed by an analysis of the QSP circuit with $U$-gates used in our algorithm. Building upon these closed-form results, we propose a phase estimation method combining Fourier analysis with QSP to separate the two gate parameters of interest in their functional forms. Our estimation algorithm enables fast and deterministic data post-processing using only direct linear algebra operations rather than iterative black-box optimizations used in multi-parameter robust phase estimation [16, 26]. Moreover, separating the inference of $\theta$ and $\varphi$ enhances the robustness of the phase estimation method against time-dependent errors that predominantly affect the gate parameter $\varphi$, which arise from the time-dependent qubit frequency noise [24, 10, 38]. The analysis and modeling of Monte Carlo sampling error also indicate that our phase estimation method achieves the fundamental quantum metrology optimality in a practical regime against realistic errors for near-term devices. We also provide a comprehensive mathematical analysis of methods based on robust phase estimation [3, 26], and prove that the vulnerability of phase angle $\varphi$ to time-dependent errors ultimately renders the estimation accuracy of the swap angle $\theta$ exponentially worse than Heisenberg limit. In addition, we proposed a noise-robust QSPE protocol that enables the estimation of gate parameters even when the gate angles fall outside the confidence regime for phase estimation. Furthermore, we demonstrated QSPE experimentally on 34 superconducting qubits using the Google Quantum AI team’s hardware, achieving $10^{-4}$ two-qubit phase estimation accuracy in practice, which improves by two orders of magnitude over standard previous methods. Lastly, we include an empirical noise-robust QSPE protocol that enables the estimation of gate parameters even when the gate angles fall outside of the approximation regime for phase estimation.
@@
Previous metrology methods [16, 26, 3] based on optimal measurements [6] for achieving the Heisenberg limit fall short of providing sufficient accuracy in $\theta$ when $\theta\ll 1$. Two significant factors limit these traditionally-regarded “optimal” metrology schemes. First, the accuracy in $\theta$ depends on the amplification factor, i.e., maximum circuit depth. The relatively low qubit coherence times of superconducting qubits render randomization-based quantum gate learning techniques [17, 21, 4] impractical due to their inefficient circuit depths. The finite low qubit coherence times [4] of superconducting qubits render randomization-based quantum gate learning techniques [17, 21, 4] impractical due to their inefficient circuit depths needed to achieve the desired accuracy close to surface code threshold [1]. Techniques based on robust phase estimation can require prohibitive depths to achieve a meaningful full signal-to-noise ratio for small $\theta$ and require iterative black-box optimizations for their estimators [26] instead of fast, deterministic post-processing for single-parameter phase estimation [16]. Second, time-dependent unitary error in $\varphi$ is prevalent in architectures like Google’s superconducting quantum computers [30], which invalidates basic assumptions in traditionally optimal and Heisenberg-limit-achieving metrology schemes.
@@
In this work, we provide a low-depth phase estimation method for estimating the angles in some invariant subspace of an unknown $U$-gate when the swap angle is small, of order below $10^{-3}$, while facing realistic time-dependent phase errors in $\varphi$. The phase estimation method leverages the structure of periodic circuits analyzed by classical and quantum signal processing and provides a framework to engineer quantum metrology from the perspective of universal quantum signal transformation. We call this type of metrology method Quantum Signal-Processing Phase Estimation (QSPE). Let $\omega$ be a tunable phase parameter and {|$\{\ket{0_{\ell}},\ket{1_{\ell}}\}$ be the logical basis of the two-level space of interest. Then, QSPE measures the transition probabilities of the quantum circuits corresponding to logical Bell states |$\ket{+_{\ell}}:=\frac{1}{\sqrt{2}}\left(\ket{0_{\ell}}+\ket{1_{\ell}}\right)$ and |$\ket{\mathrm{i}_{\ell}}:=\frac{1}{\sqrt{2}}\left(\ket{0_{\ell}}+\mathrm{i}\ket{1_{\ell}}\right)$. The transition is measured with respect to the logical basis state |$\ket{0_{\ell}}$. We depict the quantum circuit for QSPE in Figure 1 with an exemplified two-qubit $U$-gate for simplicity. In the example, the two-level subspace is set to the single-excitation subspace with basis |$\ket{0_{\ell}}=\ket{01}$ and |$\ket{1_{\ell}}=\ket{10}$. Then, the logical Bell state coincides with the conventional Bell state. We remark that the quantum circuit for QSPE can be generalized to multi-qubit cases following the recipe outlined in this paragraph. Details and the quantum circuit for QSPE in a general setup are provided in Supplementary Note 2. The transition probability corresponding to the logical Bell state $\ket{+_{\ell}}$ is denoted as $p_{X}(\omega;\theta,\varphi,\chi)$, and that corresponding to the logical Bell state |$\ket{\mathrm{i}_{\ell}}$ is denoted as $p_{Y}(\omega;\theta,\varphi,\chi)$.
@@
As outlined in Supplementary Note 3, the reconstructed function derived from measurement probabilities admits an approximated expansion $\mathfrak{h}(\omega;\theta,\varphi,\chi)=\sum_{-d+1}^{d-1}c_{k}e^{2\mathrm{i}k\omega}$ and when $d\theta\leq 1$, the coefficients are
@@
Due to Fourier expansion, sampling the reconstructed function on $(2d-1)$ distinct $\omega$ points is sufficient to characterize its information completely. For efficient processing with the Fast Fourier Transformation (FFT), we choose a set of $\omega$ points that are equally spaced. This choice of equally-spaced sampling points not only ensures numerical stability, as demonstrated in textbooks on numerical analysis [35], but also simplifies error analysis, as described in Supplementary Note 4. The second important consequence of this result is that the dependencies on $\theta$ and $\varphi$ are completely separated into the amplitude and the phase of the Fourier coefficients, respectively. The estimation problems of $\theta$ and $\varphi$ are then reduced to two independent linear regression problems. As $\chi$ is not considered, we apply a sequential phase difference to distill the angle $\varphi$:
$\bm{\Delta}=(\Delta_{0},\cdots,\Delta_{d-2})^{\top},\Delta_{k}:=\mathsf{phase}(c_{k}\overline{c_{k+1}})=2\varphi.$ (8)
Considering the Monte Carlo sampling error due to the finite number of measurements, we derive in Supplementary Note 4 the linear-regression-based estimators of the relevant angles:
$\hat{\theta}=\frac{1}{d}\sum_{k=0}^{d-1}\absolutevalue{c_{k}}\quad\text{ and }\quad\hat{\varphi}=\frac{1}{2}\frac{\bm{\bm{1}}^{\top}\mathfrak{D}^{-1}\bm{\Delta}}{\bm{\bm{1}}^{\top}\mathfrak{D}^{-1}\bm{\bm{1}}}.$ (9)
Here, $\bm{\bm{1}}$ is an all-one vector and $\mathfrak{D}$ is a $(d-1)$-by-$(d-1)$ constant tridiagonal matrix which coincides with the discrete Laplacian of a central finite difference form (see Definition 15 in Supplementary Note 4 for more details). The structure of $\mathfrak{D}$ comes from differentiating experimental noises when applying sequential phase difference. To obtain $\hat{\chi}$, we defer the task to the metrology circuit in [3, Fig. S5] and do not use QSPE for the task. The main workflow of the QSPE is depicted in Figure 2 and the Algorithm displayed in Box 1.
@@
The performance of the statistical estimators is measured by their biases and variances. In Supplementary Note 4 B, we derive the performance of QSPE estimators with the following theorem by treating QSPE as linear statistical models. Furthermore, in Supplementary Note 6, we show that QSPE estimators in Equation 9 are optimal by saturating the Cramér-Rao lower bound (CRLB) of the estimation problem.
@@
We note that the unbiasedness of these estimators holds up to a high-order bias, which is negligible in the target regime. For further details, please refer to Supplementary Note 3.
@@
According to the framework developed in Ref. [12], the variance of any quantum metrology is lower bound by the Heisenberg limit. It indicates that in our experimental setup when $d$ is large enough, optimal variance scales as $1/(d^{3}M)$. This seemingly contradicts Theorem 2, where the variance of QSPE $\varphi$-estimator can achieve $1/(d^{4}M)$. This counterintuitive conclusion is due to the pre-asymptotic regime $d\ll 1/\theta$. In Supplementary Note 6, we analyze the CRLB of QSPE. The optimal variance given by CRLB is exactly solvable in the pre-asymptotic regime $d\ll 1/\theta$.
The key reason behind such faster than Heisenberg limit scaling in pre-asymptotic regime depends on the unique structure of the QSPE circuit: the measurement outcome (Supplementary Equation (7)) concentrates around a constant value regardless of the gate parameter values. Yet when $d$ is large enough to pass to the asymptotic regime, measurement probabilities will take arbitrary values. Furthermore, the analysis of the CRLB suggests that the optimal asymptotic variance agrees with the Heisenberg limit when $d$ is large enough. This nontrivial transition of optimal variance is theoretically analyzed and numerically justified in Supplementary Note 6. We summarize this nontrivial transition of the optimal variance scaling of QSPE as a phase diagram in Figure 3 a. To numerically demonstrate the transition, we compute the exact CRLB of QSPE when $\theta=1\times 10^{-2}$ and $\theta=1\times 10^{-3}$. In Figure 3 b, the slope of the curve in log-log scale exhibits a clear transition before and after $d=1/\theta$, which supports the phase diagram in Figure 3 a. Detailed theoretical and numerical discussions of the transition are carried out in Supplementary Note 6.
@@
Analyzing Cramér-Rao bounds suggests the optimality of a quantum metrology protocol or the suboptimality leading to further improvement. Given an initialization and measurement, the optimality lies in the analysis of the classical CRLB, which investigates the most information one can retrieve from measurement probabilities. As outlined in Supplementary Note 6 A, the CRLBs are solvable when $d\ll 1/\theta$, which exactly agree with the variance of our estimators derived in Theorem 2. The optimality of our estimator is also validated from numerical simulation depicted in Figure 3 b. Such agreement reveals the optimality of our data post-processing. Although linear-regression-based estimators are used, this linearization does not sacrifice the information retrieval in the experimental data. Furthermore, in contrast to other iterative inference methods, our estimators directly estimate angles using basic linear algebra operations, to which stability and fast processing are credited.
Despite the informative indication by analyzing CRLB, it cannot provide direct suggestions on improving initialization and measurement. Such generalization demands the switch to quantum Cramér-Rao lower bound (QCRLB) which requires upper bounding the quantum Fisher information (QFI). As a quantum analog of the classical Fisher information, QFI lies in the center of quantum metrology by providing a fundamental lower bound on the accuracy one can infer from the system of a given resource limit. According to the analysis in [18], the QFI is an upper bound on the Fisher information over all possible measurements. For brevity, we only consider the inference of $\theta$ and hold all other unknown parameters constant in the analysis. However, our analysis can be generalized to the multiple parameter inference by adopting the multi-variable QFI in Ref. [18]. In Supplementary Note 6 D, we derive that the average QFI is upper bounded by an integral:
@@
Here, the integrand gauges the information contained in an experiment with angle $\omega$. The integral stands for the use of equally spaced $\omega$ samples due to the absence of accurate information of $\varphi$. It is worth noting that the integrand is sharply peaked at $4d^{2}$ when $\omega=\varphi$ which is also referred to as the phase-matching condition. The missing information of $\varphi$ downgrades the average QFI from $4d^{2}$ to $4d$. However, as discussed in the following subsection, the lack of information about $\varphi$ in existing methods can potentially significantly degrade the Fisher information to $\mathcal{O}(\log(d))$, thereby severely impeding the achievement of Heisenberg-limit scaling in estimation accuracy. Furthermore, this also suggests a finer estimation of $\theta$ when some rough information of $\varphi$ is provided either as a priori or from some preliminary estimation. This improvement is discussed in Supplementary Note 4 C. Consequently, the QCRLB of the QSPE formalism is
@@
Compared with Theorem 2, we see differentiation in a constant suboptimal factor of $2$, which is explainable. Note that we use two logical Bell states to perform experiments. The advantage is the experimental probabilities of these two experiments form a conjugate pair to reconstruct a complex function for the ease of analysis. This complex function and its properties (see Theorems 7 and 9 in Supplementary Note 3) eventually lead to a simple robust statistical estimator requiring only light computation. In contrast, the data generated from the initialization of one Bell state still contains full information on the parameters to be estimated. However, the highly nonlinear and oscillatory dependency renders the practical inference challenging. Hence, the factor of $2$ is due to the use of a pair of Bell states. Although the QFI indicates that inference variance can be lower by removing such redundancy in the initialization, the nature of ignoring practical ease makes it hard to achieve.